OpenStudy (anonymous):

Can someone PLEASE help me solve this equation and put it in decimal form :/ im Stuck. log2 9 + log2 a = log2 11 all the 2's are little 2's (there subscript)

OpenStudy (anonymous):

\[\log_2(9)+\log_2(a)=\log_2(9a)\] by property of the log that says \[\log(A)+\log(B)=\log(AB)\]

OpenStudy (anonymous):

then if \[\log_2(9a)=\log_2(11)\] it must be true that \[9a=11\]because any log is a one to one function

OpenStudy (anonymous):

so 1.2222 would be the answer correct?

OpenStudy (anonymous):

i would say it is \(\frac{11}{9}\) and leave it at that

OpenStudy (anonymous):

lol thanks so much